# Mental Disorders Detection in the Era of Large Language Models

### v1

@@ Abstract
This paper compares the effectiveness of traditional machine learning methods, encoder-based models, and large language models (LLMs) on the task of detecting depression and anxiety. Five datasets were considered, each differing in format and the method used to define the target pathology class. We tested AutoML models based on linguistic features, several variations of encoder-based Transformers such as BERT, and state-of-the-art LLMs as pathology classification models. The results demonstrated that LLMs outperform traditional methods, particularly on noisy and small datasets where training examples vary significantly in text length and genre. However, psycholinguistic features and encoder-based models can achieve performance comparable to language models when trained on texts from individuals with clinically confirmed depression, highlighting their potential effectiveness in targeted clinical applications.
Mental Disorders Detection in the Era of Large Language Models
Gleb Kuzmin2,4 Petr Strepetov3 Maksim Stankevich4 Ivan Smirnov4 Artem Shelmanov1 1MBZUAI 2AIRI 3MIPT 4FRC CSC RAS kuzmin@airi.net strepetov.pa@phystech.edu stankevich@isa.ru ivs@isa.ru artem.shelmanov@mbzuai.ac.ae
@@ 1 Introduction
The problem of detecting mental disorders and patient emotions through text analysis and machine learning has been of increasing interest to researchers over the past decade (Graham et al., 2019; Zhang et al., 2022; Calixto et al., 2022; Mayer et al., 2024). In fact, advances in data science and natural language processing methods offer promising opportunities for screening, monitoring, early detection, and prevention of negative outcomes of mental disorders. Although there are studies that work with interviews (Morales and Levitan, 2016; Ringeval et al., 2017) and offline texts (Lynn et al., 2018; Stankevich et al., 2019), in most cases the material for such research comes from social media (Guntuku et al., 2017; Garg, 2023). These studies tend to focus on but are not limited to, conditions such as depression, anxiety, stress, suicidality, post-traumatic stress disorder, and anorexia. Unsurprisingly, the methods considered for predicting mental state from text fell into traditional machine learning, using hand-crafted linguistic features, and various forms of deep learning (Zhang et al., 2022). The deep learning approach is often more accurate, especially when there are enough data samples, while traditional machine learning produces more interpretative results.
In this paper, we compare the performance of linguistic features, encoder-based models, and large language models (LLMs) on the task of identifying mental disorders. We consider two types of mental states, depression and anxiety, and several datasets in Russian that differ in text format and in the way a pathology is detected.
We are particularly interested in evaluating the performance on the essay dataset because it consists of texts written by patients with clinically diagnosed depression and texts written by volunteers who were considered healthy (Stankevich et al., 2019). It is worth pointing out the problem of dividing data into texts written by people with a certain mental pathology. In most cases, social media-based studies use self-report, affiliation, or questionnaire data to determine mental health status. However, there is a lack of work using clinically validated data, which can vary widely (Chancellor and De Choudhury, 2020; Ernala et al., 2019). We believe that verified training data from essays can significantly contribute to building models for the social media domain.
@@
We conducted a thorough comparison of various groups of models on the depression and anxiety detection tasks in Russian, which could be used for the practitioners in this field for future experiments.
We investigated the transferability of models from the task with clinical diagnosis as a target to the task with a target obtained as a questionnaire result to alleviate the lack of clinically validated data in the mental disorder detection task.
@@ 2.1 Traditional and Advanced ML Methods
Researchers have employed diverse methodologies and techniques in a comprehensive exploration of depression and anxiety detection across various social media platforms. Tadesse et al. (2019) consider the Reddit users’ dataset, comparing single and combined feature learning for depression detection. N-gram features, Linguistic Inquiry and Word Count (LIWC) dictionary features, and topics from Latent Dirichlet Allocation are considered, showcasing the effectiveness of combined features on the multi-layer perceptron with a remarkable 91% accuracy and 0.93 F1 score.
In (Shah et al., 2020), the authors leverage NLP methods for depression detection of Reddit users based on their posts. GloVe, Word2Vec, and FastText embeddings, as well as handcrafted statistical metadata features and features from the LIWC dictionary, are used for text representation. The two-headed model, combining BiLSTM for embedding transformation and a fully connected layer for meta-features, demonstrates superior results with Word2Vec embeddings and meta-features achieving a noteworthy 0.81 F1 score. Additionally, the authors use Early Risk Detection Error and Latency metrics to take the time of classification into account.
The study (Owen et al., 2020) considers depression and anxiety detection for Twitter posts. SVM on TF-IDF vectors and GloVe embeddings, along with BERT (Devlin et al., 2019) and ALBERT (Lan et al., 2019), are explored, showcasing the nuanced performance of BERT on a balanced dataset and better performance of SVM on an unbalanced one.
In a comprehensive survey (Babu and Kanaga, 2021), the importance of emoticons in texts in sentiment analysis for depression detection is underscored. The study covers 101 publications, emphasizing the effectiveness of combining deep learning algorithms, with CNNLSTM yielding the highest precision. Moreover, multi-class classification, categorizing sentiments into subclasses based on polarities, provides more precise results than binary and ternary classification.
In (Tejaswini et al., 2022), the authors introduce the FCL (FasttextCNNLSTM) model, comparing it with LSTM and CNN models on Glove and Word2vec embeddings. They also use datasets containing Reddit and Twitter posts. FCL outperforms other models in identifying depression in both small and large datasets.
Assessing the anxious Twitter posts caused by the COVID-19 pandemic, (Jeong et al., 2023) delves into anxiety detection using BERT trained on the Korean language, achieving a 0.67 macro F1-score and 0.91 accuracy and establishing a correlation between the anxiety index and COVID-19 waves.
The study (Ansari et al., 2023) focuses on identifying depression in social media datasets (CLpsych, Reddit, eRisk) through various text classification methods, combining sentiment lexicons with deep learning pipelines. Authors utilize sentiment lexicons with logistic regression and LSTM with attention on Glove embeddings, comparing them and combining them into ensembles. Ensemble models outperform hybrid lexicon- and DL-based models on all datasets, achieving 75% accuracy and 0.77 F1 score on the Reddit dataset.
@@ 2.2 Large Language Models
According to the systematic review (Omar and Levkovich, 2024), most studies about depression detection focus on BERT-based models, indicating the field’s early stages in adopting newer technologies like GPT-4 and Google’s Gemini. However, LLMs are demonstrating significant potential in enhancing the capabilities of depression detection systems.
The Chat-Diagnose approach (Qin et al., 2023) integrates diagnostic criteria from the Diagnostic and Statistical Manual of Mental Disorders (DSM-5) into prompts and uses the Chain of Thoughts technique to deliver explainable diagnoses via an LLMs-augmented system based on ChatGPT/GPT-3. This method demonstrates state-of-the-art results on Twitter and Weibo depression datasets by employing zero-shot and few-shot learning, using prompts with selected social media content alongside predictions from a Time2VecTransformer model to enhance diagnostic accuracy.
@@
Additionally, research (Wang et al., 2024) on eRisk 2023 and eRisk 2021 datasets explores detecting depression symptoms and classifying depression severity using LLMs. Utilizing Beck’s Depression Inventory to form queries related to depression symptoms and the Universal Sentence Encoder for text embeddings, the study creates two datasets containing top-1 and top-5 ranked texts for each query. LLMs fine-tuned with QLoRA, including Llama-2-13b-chat, SUS-Chat-34B, and Neural-chat-7b-v3, are used for classification into four levels of depression severity. The top-five dataset performs better overall, with the combined model neural-chat+SUS-Chat_top5 excelling in Average Hit Rate and Average Closeness Rate, while neural-chat_top5 and SUS-Chat_top5 achieve the highest Average DODL and Depression Category Hit Rate among single models.
The DORIS (Lan et al., 2024) system addresses the challenges of detecting depression through social media posts from the Sina Weibo Depression Dataset. The authors use GPT3.5-Turbo-1103 for annotating high-risk texts according to the DSM-5 depression scale, also LLM is used to summarize critical information from users’ historical mood records (mood courses). The final model based on XGBoost is learned on features from annotations and gte-small-zh model vector representations of post histories and mood courses. It shows an improvement in the AUPRC metric by 0.036 compared to baseline methods. This study underscores the system’s potential for early detection and intervention, providing a valuable tool for mental health assessment on social media platforms.
We are the first to examine and compare three generations of the discussed models for depression and anxiety detection tasks in Russian. Unlike other works, we used different models from each group and carefully compared the results of the models between the groups on five datasets, aiming for a general recommendation on the best models to use in practice.
@@ 3 Data
Classes in all datasets were represented in the binary format: a healthy class and a pathology class. The general description of the datasets used in our study is shown in Table 1.
| Label | DE | DSM | AL | AD | AC | ||
@@ 3.1 Depression-Essays
To compile this dataset, subjects were asked to write a short essay (up to 5,000 characters) on the topic of “Myself, others, world” (Stankevich et al., 2019). A total of 557 essays were collected, where 110 of them were written by people with clinical depression. The essays written by people with clinically validated depression were provided by the Mental Health Research Center, Moscow, Russia. The label DE (Depression-Essay) is further used to describe this dataset. The best-reported performance on this data reaches 73% F1 for the depression class in the cross-validation evaluation with the random forest model trained on n-grams and psycholinguistic features.
3.2 Depression-Social Media
@@
The other depression dataset under consideration was created from data collected to study the problem of identifying depression in social network users (Ignatiev et al., 2022). The best-reported classification performance using textual data reaches only 65% F1 for the depression class with a logistic regression model trained on psycholinguistic features. This dataset contains text messages from the social network VKontakte and results from the Beck Depression Inventory (BDI) questionnaire (Beck et al., 1987). The healthy class included users with a scale score of 10 or less, and the pathological class between 30 and 63.
@@
We have made some changes to the original dataset. For each subject, all messages were combined for a period of 170 days prior to the date of the questionnaire screening, and the total text was limited to 6,000 characters. Such restrictions were imposed to bring the final texts from this dataset closer to the format of the essay dataset in terms of total text length for each subject and to account for the fact that the relevance of depression screening becomes less over a period of more than six months. The label DSM (Depression-Social Media) is used hereafter to describe this dataset.
3.3 Anxiety-Letter and Picture Description
@@
The RusNeuroPsych corpus (Litvinova and Ryzhkova, 2018) was used as the data for creating the anxiety detection model. This corpus was prepared to study the relationships between a person’s text, personal traits, mental status, and demographic characteristics. To compile it, subjects were asked to write an informal letter to a friend and give a textual description of a picture, and then complete a series of psychological questionnaires. Among the questionnaires, the Hospital Anxiety and Depression Scale (HADS) (Bjelland et al., 2002) was used. Thus, the data were divided using the results of this questionnaire, where the healthy class included subjects with a score of 7 and below, and the pathology class included all subjects with a score of 8 and above. Since the corpus contained 2 different types of text, it was divided into 2 separate datasets for the experiments. It is worth noting that the results of the HADS questionnaire also allow us to divide subjects into groups based on the presence and absence of depression, but a number of subjects with high depression scale scores was less than 20, which prevented us from conducting experiments to train classification models. The labels AL (Anxiety-Letter) for letters and AD (Anxiety-Description) for picture description are further used to describe these datasets. To the best of our knowledge, no classification experiments have been performed on this data before.
3.4 Anxiety-COVID Comments
@@
In addition, we used a corpus of subjects’ comments on the COVID-19 pandemic situation (Medvedeva et al., 2021). Subjects were asked to complete a series of questionnaires and write a free-form commentary describing their attitudes towards the world situation around the pandemic and self-isolation. Among the questionnaires used was the SCL-90-R symptom questionnaire (Derogatis, 1983), the anxiety scale from which was used to form 2 groups: the healthy group, with an anxiety scale score below the 33rd percentile, and the pathology group, with an anxiety scale score above the 66th percentile. The label AC (Anxiety-Covid) is used hereafter to describe this dataset. To the best of our knowledge, no classification experiments have been performed on this data before.
4 Method
@@ 4.1.1 Psycholinguistic Features
The linguistic features used in our research were calculated using the tool described in (Smirnov et al., 2021). This tool allows the calculation of morphological, syntactic, and vocabulary parameters of texts, including various psycholinguistic coefficients. A total of 113 features were used. A detailed description of the features used and the features obtained with this tool on our data is given in the Huggingface repository111https://huggingface.co/datasets/anonymizedauthor/paper_data.
@@ 4.1.2 Classification Setup
As a classification baseline, we use AutoML system auto-sklearn (Feurer et al., 2015) and learn it on psycholinguistic features and on n-grams. Auto-sklearn classifier employs a Bayesian optimizer which considers 15 classification algorithms, 14 feature preprocessing methods, and 4 data preprocessing methods. Also, the classifier utilizes a meta-learning approach and automated ensemble construction to speed up optimization.
For psycholinguistic features, we consider several methods of feature selection: filter method, wrapper method (forward selection and backward elimination), and embedded method. Selected features are examined for linear correlation and ones with absolute Pearson coefficient values more than 0.95 are deleted.
We also train models on TF-IDF vectors on unigrams and unigrams with bigrams. In addition to full vectors, we consider different subsets of features for vectors on unigrams: from 20% to 100% of the features are selected with 20% increment. For each subset, correlated features are deleted in the same way as with psycholinguistic features.
On each set of features we launch the classifier 6 times with different seeds and then calculate the mean and standard deviation values of the metrics based on the results of the launches.
4.2 Encoder Models
We employ classification with encoder models to compare methods, based on psycholinguistic features, with more modern solutions. As we target classification on Russian corpora, we consider models pretrained on multilingual or Russian datasets. As a simple baseline, we used a base version of multilingual BERT, a model with 110 million parameters, first introduced in (Devlin et al., 2019). Another considered baseline is RuBERT (Kuratov and Arkhipov, 2019), a pretrained Russian version of BERT-base with the same amount of parameters. We also finetuned several more recent models, such as RuBioRoBERTa (Yalunin et al., 2022) which is a RoBERTa, pretrained on Russian language biomedical texts, and RuRoberta-large - a bigger version of RoBERTa, also pretrained on Russian language datasets. For all of the four models, we optimized hyperparameters using Bayesian optimization from HuggingFace (Wolf et al., 2019) framework.
The used hyperparameter grid and optimal parameters, alongside with used checkpoint of models, are presented in Appendix A.

### v3

# Exploring Large Language Models for Detecting Mental Disorders

@@ Abstract
This paper compares the effectiveness of traditional machine learning methods, encoder-based models, and large language models (LLMs) on the task of detecting depression and anxiety. Five Russian-language datasets were considered, each differing in format and in the method used to define the target pathology class. We tested AutoML models based on linguistic features, several variations of encoder-based Transformers such as BERT, and state-of-the-art LLMs as pathology classification models. The results demonstrated that LLMs outperform traditional methods, particularly on noisy and small datasets where training examples vary significantly in text length and genre. However, psycholinguistic features and encoder-based models can achieve performance comparable to language models when trained on texts from individuals with clinically confirmed depression, highlighting their potential effectiveness in targeted clinical applications.111https://github.com/glkuzi/llm-mental-disorders-detection
Exploring Large Language Models for Detecting Mental Disorders
Gleb Kuzmin1,2,3 Petr Strepetov4 Maksim Stankevich3 Natalia Chudova3 Artem Shelmanov6 Ivan Smirnov3,5 1AIRI 2ISP RAS Research Center for Trusted Artificial Intelligence 3FRC CSC RAS 4MIPT 5RUDN University 6MBZUAI kuzmin@airi.net strepetov.pa@phystech.edu stankevich@isa.ru nchudova@gmail.com artem.shelmanov@mbzuai.ac.ae ivs@isa.ru
@@ 1 Introduction
The problem of detecting mental disorders and patient emotions through text analysis and machine learning has been of increasing interest to researchers over the past decade (Graham et al., 2019; Zhang et al., 2022; Calixto et al., 2022; Mayer et al., 2024). In fact, advances in data science and natural language processing methods offer promising opportunities for screening, monitoring, early detection, and prevention of negative outcomes of mental disorders. Although there are studies that work with interviews (Morales and Levitan, 2016; Ringeval et al., 2017) and offline texts (Lynn et al., 2018; Stankevich et al., 2019), in most cases the material for such research comes from social media (Guntuku et al., 2017; Garg, 2023). These studies tend to focus on, but are not limited to, conditions such as depression, anxiety, stress, suicidality, post-traumatic stress disorder, and anorexia. Unsurprisingly, the methods considered for predicting mental state from text fell into traditional machine learning, using hand-crafted linguistic features, and various forms of deep learning (Zhang et al., 2022). The deep learning approach is often more accurate, especially when there are enough data samples, while traditional machine learning produces more interpretable results.
In this paper, we compare the performance of linguistic features, encoder-based models, and large language models (LLMs) on the task of identifying mental disorders. We consider two types of mental states, depression and anxiety, and several datasets in Russian that differ in text format and in the way a pathology is detected. To the best of our knowledge, this is the first comprehensive study on mental disorder detection in Russian texts, involving a wide range of techniques, including LLMs.
For our study, we involve a clinically verified dataset of essays, which contains texts written by patients with clinically diagnosed depression as well as texts from healthy volunteers (Stankevich et al., 2019). Most studies based on social media rely on self-reports, group affiliations, or questionnaire responses to determine mental health status. In contrast, only a few works use clinically validated data, which can differ substantially in quality and reliability (Chancellor and De Choudhury, 2020; Ernala et al., 2019).
@@
RQ3: How do the LLM-generated explanations for detected depression align with the clinician’s perspective?
@@
We conducted a thorough comparison of various groups of models on the depression and anxiety detection tasks in Russian, which could be used by practitioners in this field for future experiments.
We explored the transferability of models from tasks using clinical diagnoses as targets to those based on questionnaire-derived labels, aiming to mitigate the scarcity of clinically validated data in mental disorder detection.
We evaluated and categorized LLM-generated explanations for detected depression from the point of view of clinicians, which could be used for future improvement of LLM-assisted systems in this field.
@@ 2.1 Traditional and Advanced ML Methods
Researchers have employed various methods for detecting depression and anxiety across social media platforms. Tadesse et al. (2019) consider the Reddit users’ dataset, comparing single and combined feature learning for depression detection. N-gram features, Linguistic Inquiry and Word Count (LIWC) dictionary features, and topics from Latent Dirichlet Allocation are considered, showcasing the effectiveness of combined features on the multi-layer perceptron.
In (Shah et al., 2020), NLP methods are applied for depression detection of Reddit users based on their posts. GloVe, Word2Vec, and FastText embeddings, as well as handcrafted statistical metadata features and LIWC features, are used for text representation. The two-headed model, combining BiLSTM for embeddings and a fully connected layer for meta-features, demonstrates superior results with Word2Vec embeddings and meta-features. Additionally, the authors use Early Risk Detection Error and Latency metrics to take the time of classification into account.
Owen et al. (2020) consider depression and anxiety detection for tweets, using SVM on TF-IDF vectors and GloVe embeddings, as well as BERT (Devlin et al., 2019) and ALBERT (Lan et al., 2020). Results show that BERT is better on a balanced dataset, while SVM excels on an unbalanced one.
Babu and Kanaga (2022) underscores the importance of emoticons in texts in sentiment analysis for depression detection. The study covers 101 publications, emphasizing the effectiveness of combining deep learning algorithms, with CNNLSTM yielding the highest precision. Moreover, multi-class sentiment classification provides more precise results than binary and ternary classifications.
The FCL (FasttextCNNLSTM) model proposed in (Tejaswini et al., 2024) outperforms LSTM and CNN models, based on GloVe and Word2vec embeddings, on datasets for depression detection with Reddit and Twitter posts.
Assessing the anxious Twitter posts caused by the COVID-19 pandemic, (Jeong et al., 2023) uses BERT trained on the Korean language, achieving strong accuracy and establishing a correlation between the anxiety index and COVID-19 waves.
The study (Ansari et al., 2023) focuses on identifying depression in social media datasets (CLPsych, Reddit, eRisk) through various text classification methods, combining sentiment lexicons with deep learning pipelines. Authors utilize sentiment lexicons with logistic regression and LSTM with attention on GloVe embeddings, comparing them and combining them into ensembles.
@@ 2.2 Large Language Models
According to the systematic review (Omar and Levkovich, 2025), most studies about depression detection focus on BERT-based models, indicating the field’s early stages in adopting newer technologies like GPT-4 and Google’s Gemini. However, LLMs are demonstrating significant potential in advancing depression detection systems.
The Chat-Diagnose approach (Qin et al., 2023) integrates diagnostic criteria from the Diagnostic and Statistical Manual of Mental Disorders (DSM-5) into prompts and uses the Chain of Thoughts technique to deliver explainable diagnoses via an LLMs-augmented system based on ChatGPT/GPT-3. This method demonstrates state-of-the-art results on Twitter and Weibo depression datasets by employing zero-shot and few-shot learning.
@@
Additionally, Wang et al. (2024) investigates depression symptom detection and severity classification using LLMs on the eRisk 2021 and eRisk 2023 datasets. Utilizing Beck’s Depression Inventory to form queries related to depression symptoms and the Universal Sentence Encoder for text embeddings, the study creates two datasets containing top-1 and top-5 ranked texts for each query. LLMs fine-tuned with QLoRA are used for classification into four levels of depression severity.
The DORIS (Lan et al., 2024) system addresses the challenges of detecting depression through social media posts from the Sina Weibo Depression Dataset. The authors use GPT3.5-Turbo-1103 for annotating high-risk texts according to the DSM-5 depression scale; also, LLM is used to summarize critical information from users’ historical mood records (mood courses). The final model based on XGBoost is learned on features from annotations and gte-small-zh model vector representations of post histories and mood courses and shows an improvement over the baseline.
We are the first to examine and compare three generations of the discussed models for depression and anxiety detection tasks in Russian, namely, traditional ML models, encoder-decoder models, and LLMs. Unlike other works, we used various models from each group and carefully compared the results of the models between the groups on five datasets, aiming for a general recommendation on the best models to use in practice.
@@ 3 Data
This paper considers five Russian-language datasets: 2 for depression and 3 for anxiety. Classes in all datasets were represented in the binary format: a healthy class (no signs of mental disorders) and a pathology class (depression or anxiety). The general description of the datasets used in our study is shown in Table 1.
| Name | Depression-Essays (DE) | Depression-Social Media (DSM) | Anxiety-Letter (AL) | Anxiety-Description (AD) | Anxiety-COVID Comments (AC) | ||
@@ 3.1 Depression-Essays (DE)
To compile this dataset, subjects were asked to write a short essay (from 1,500 to 5,000 characters with spaces and punctuation) on the topic of “Myself, others, world” (Stankevich et al., 2019). In total, 557 essays were collected, including 110 authored by patients with clinical depression. The essays written by people with clinically validated depression were provided by the Mental Health Research Center, Moscow, Russia. The collection of essays was done on a voluntary basis, under conditions of anonymity and for research purposes only. The best-reported performance on this data reaches 73% F1 for the depression class in the cross-validation evaluation with the random forest model trained on n-grams and psycholinguistic features (Stankevich et al., 2019). For this dataset, we provided several anonymized examples in Table˜21 in Appendix˜G.
3.2 Depression-Social Media (DSM)
@@
The Depression-Social Media dataset was developed to support research on detecting depression in social network users (Ignatiev et al., 2022). It consists of text messages from the VKontakte platform, accompanied by results from the Russian-language adaptation of the 21-item Beck Depression Inventory (BDI) questionnaire (Beck et al., 1961). The healthy class included users with a scale score of 10 or less, and the pathological class between 30 and 63. The best-reported classification performance using textual data reaches only 65% F1 for the depression class with a logistic regression model trained on psycholinguistic features (Ignatiev et al., 2022).
@@
We have made some changes to the original dataset. For each subject, all messages were combined for a period of 170 days prior to the date of the questionnaire screening, and the total text was limited to 6,000 characters (symbols), resulting in a median length per sample of 122 words. Such restrictions were imposed to bring the final texts from this dataset closer to the format of the essay dataset in terms of total text length for each subject and to account for the fact that the relevance of depression screening becomes less over a period of more than six months.
3.3 Anxiety-Letter and Anxiety Picture Description (AL and AD)
@@
For anxiety detection experiments, we use the RusNeuroPsych corpus (Litvinova and Ryzhkova, 2018). This corpus was prepared to study the relationships between a person’s text, personal traits, mental status, and demographic characteristics. To compile the corpus, participants were asked to write an informal letter to a friend, provide a textual description of a picture, and complete a series of psychological questionnaires. Among these, the Hospital Anxiety and Depression Scale (HADS) (Bjelland et al., 2002) was employed to assess anxiety levels. Based on HADS scores, subjects with a score of 7 or below were assigned to the healthy class, while those with a score of 8 or higher were assigned to the pathology class. Because the corpus contains two distinct text types, it was further split into two separate datasets for the experiments. We denote them AL (Anxiety-Letter) for letters and AD (Anxiety-Description) for picture descriptions. To the best of our knowledge, no classification experiments have been performed on this data before.
3.4 Anxiety-COVID Comments (AC)
@@
In addition, we used a corpus of subjects’ comments on the COVID-19 pandemic situation (Medvedeva et al., 2021). Subjects were asked to complete a series of questionnaires and write a free-form commentary describing their attitudes towards the world situation around the pandemic and self-isolation. Among the questionnaires used was the SCL-90-R symptom questionnaire (Derogatis, 1983), the anxiety scale from which was used to form 2 groups: the healthy group, with an anxiety scale score below the 33rd percentile, and the pathology group, with an anxiety scale score above the 66th percentile. To the best of our knowledge, no classification experiments have been performed on this data before.
4 Methods
@@ 4.1.1 Psycholinguistic Features
The linguistic features used in our research were extracted using the tool described in (Smirnov et al., 2021). This tool extracts morphological, syntactic, and vocabulary parameters of texts, including various psycholinguistic coefficients. A total of 113 features were used. A detailed description of the features utilized and those extracted with this tool on our data is available in the Hugging Face repository.222https://huggingface.co/datasets/anonymizedauthor/paper_data
@@ 4.1.2 Classification Setup
As a classification baseline, we use the AutoML system auto-sklearn (Feurer et al., 2015) trained on psycholinguistic features and n-grams. The auto-sklearn classifier employs a Bayesian optimizer that considers 15 classification algorithms, 14 feature preprocessing methods, and 4 data preprocessing techniques. Additionally, the classifier utilizes a meta-learning approach and automated ensemble construction to speed up optimization.
For psycholinguistic features, we consider several feature selection methods: filter method, wrapper method (forward selection and backward elimination), and embedding method. Selected features are examined for pairwise linear correlation, and those with absolute Pearson coefficient values of more than 0.95 are deleted.
We also train models on TF-IDF vectors on unigrams and unigrams with bigrams. In addition to full vectors, we consider different subsets of features for vectors on unigrams: from 20% to 100% of the features are selected with a 20% increment. For each subset, correlated features are deleted in the same way as with psycholinguistic features.
On each set of features, we launch the classifier 6 times with different seeds and then calculate the mean and standard deviation values of the metrics based on the results of the launches. The chosen AutoML models are presented in Appendix˜A.
4.2 Encoder-Based Models
Encoder-based classification models eliminate the need for tedious feature engineering in favor of a deep multi-layer neural architecture. As we target classification on Russian corpora, we consider models pretrained on multilingual or Russian datasets. As a simple baseline, we used a base version of multilingual BERT, a model with 110 million parameters, first introduced in (Devlin et al., 2019). Another considered baseline is RuBERT (Kuratov and Arkhipov, 2019), a pretrained Russian version of BERT-base with the same amount of parameters. We also finetuned several more recent models, such as RuBioRoBERTa (Yalunin et al., 2022), which is a RoBERTa pretrained on Russian language biomedical texts, and RuRoBERTa-large – a bigger version of RoBERTa, also pretrained on Russian language datasets. For all of the four models, we optimized hyperparameters using Bayesian optimization from the HuggingFace framework (Wolf et al., 2020).
The used hyperparameter grid and optimal parameters, along with the used checkpoint of models, are presented in Appendix˜A.